# MusicInfuser: Making Video Diffusion Listen and Dance

### v1

@@ 4.3 Beta-Uniform Scheduling
4.4 Utilizing In-the-Wild Videos
Training exclusively on highly structured datasets [46, 32] can lead to reduced generalizability and model degradation when confronted with diverse real-world scenarios. Therefore, we use a mixture of in-the-wild data with structured datasets. We gathered the data from YouTube playlists showing dance performances where we extracted video clips for training. These videos introduce valuable diversity in terms of camera angles, lighting conditions, performance environments, and dance styles. The inclusion of in-the-wild data serves as regularization, preventing overfitting to specific dance patterns or environmental settings.
4.5 Obtaining and Diversifying Captions
Furthermore, we randomly replace a small portion of detailed captions with basic, simple captions. This way, the adapter network learns how to respond to the music without relying on the text. This effectively reduces the model’s dependence on textual cues and encourages it to develop stronger associations between musical features while still maintaining prompt adherence. See the supplementary material for the exact prompt templates we use for the diversification and replacement.
4.6 Selecting Layers Based on Adaptability
Applying cross-attention mechanisms to all layers of the model is computationally expensive and may introduce unnecessary complexity. However, exhaustively searching for an optimal combination of layers is intractable. To address this, we devise a novel selection criterion to selectively apply audio conditioning based on layer adaptability. Our method draws inspiration from Spatiotemporal Skip Guidance (STG) [26], which shares our motivation of identifying layers where modulation influences motion and structure without significantly causing predictions to deviate from the originally learned manifold. See the supplementary material for details on the selected layers.
Figure 6: Speed control. The audio input is slowed down (the top row) or sped up (the bottom row) by 0.75 times and 1.25 times, respectively. This shows that speeding up generally results in more movements. Also see the change in the dynamicity, as speeding up the audio also increases the tone of the music.
5 Experiments
5.1 Implementation Details
Dataset
The AIST dataset [46] includes 13,940 videos with 60 musical pieces, 10 dance genres, and 35 dancers. We utilize AIST as one of our primary data sources for training our audio-driven dance generation model. We extract 2,378 clips and strictly divide training and test sets with non-overlapping music tracks, following [32]. For each training instance, we randomly sample approximately 2.5-second clips from the full sequences. To enhance the quality and generalizability of our model, we supplement the structured AIST data with in-the-wild dance videos. Specifically, we extract 15,799 video clips from YouTube dance videos, encompassing a wide variety of dance styles, settings, and production qualities. These clips are mixed with those from AIST at a 1:1 ratio during training, creating a balanced dataset that leverages both the controlled environment of AIST and the rich diversity of real-world dance performances.
Model Details
We train our model on a single NVIDIA A100 GPU for 4,000 steps with a learning rate of $1\text{e}{-4}$, which takes roughly 20 hours to complete. Our HR-LoRA uses rank 64, providing sufficient capacity to capture complex dance movements while maintaining parameter efficiency. For Beta-Uniform scheduling, we set initial $\beta=3$ with exponential decay toward $\beta=1$. We use Mochi [44] as our base model with a classifier-free guidance scale of $\gamma_{\text{cfg}}=6.0$ during inference and employ Wav2Vec 2.0 [5] as the audio encoder.
Quantitative Metrics
Evaluating generated content with plausible metrics presents a significant challenge. Inspired by VBench’s use of Visual-Language Models (VLMs) [31] for text-to-video assessment, we propose a novel metric based on Video-LLMs [14] to evaluate generated videos alongside their conditioning audio and prompts. Specifically, we leverage VideoLLaMA 2 [14] and formulate targeted queries to assess three key components: dance quality, video quality, and prompt alignment. For dance quality, we evaluate style alignment, beat alignment, body representation, movement realism, and choreography complexity. Video quality assessment includes imaging quality, aesthetic quality, and overall consistency. Prompt alignment evaluation covers style capture, creative interpretation, and overall satisfaction. In Table 1 and 2, we present results on the AIST test data to demonstrate the plausibility. In dance quality metrics, the AIST test data—which should serve as an upper bound for some metrics such as beat alignment, imaging quality, and movement realism—outperforms other models (including ours) in such metrics.
@@ 5.1 Implementation Details
Table 4: Ablation study. Feature addition denotes that we spatially expand the audio feature and add the feature to the corresponding frame.
5.2 Experimental Results
Music- and Text-Driven Dance Video Generation
Fig. 1 showcases the model’s ability to combine textual control with musical synchronization. The generated videos successfully incorporate scene contexts (restaurant kitchen, beach at sunset) and dancer attributes (wearing a leather jacket, chef’s uniform) as specified in the prompts, while simultaneously aligning the choreographic style with the musical input. Fig. 2 demonstrates that our model can generalize to generate group dance videos by simply modifying the prompt to reference multiple dancers.
Music Responsiveness
Figure 7: Videos generated with three distinct in-the-wild music tracks created with SUNO AI. For each row, we use in-the-wild music tracks generated with a word “K-pop,” an unseen category.
Generalization to In-the-Wild Music and Longer Videos
To evaluate generalization beyond the AIST music distribution, we test our model with music tracks generated by SUNO AI, which represent styles not represented in the training data. Fig. 7 shows successful generation for these unseen music categories, confirming the model’s ability to map novel audio patterns to appropriate dance movements. Fig. 3 shows longer video generation results with the same setting but with twice as many frames as the videos we used for training. These experiments show flexible generalization of our method.
Comparison with Prior Work
Fig. 5 presents a side-by-side comparison with MM-Diffusion [39]. MusicInfuser produces more consistent human forms, fewer visual artifacts, and more fluid, realistic movements compared to both baselines. Notably, our method also demonstrates superior alignment between music beat patterns and dance movements, visible in the correspondence between spectrogram intensity and motion transitions.
Our approach overcomes key limitations of prior work: unlike MM-Diffusion, which generates shorter videos with limited style control, MusicInfuser produces longer sequences with both musical synchronization and prompt-based style control, while improving overall consistency of the video and having fewer artifacts. Compared to Mochi, our method adds music responsiveness while maintaining or improving video quality and prompt adherence.
Quantitative Evaluation
Tables 12 present our quantitative results compared against several baselines [39, 44]. For dance quality (Table 1), our method outperforms previous approaches in style alignment, beat alignment, movement realism, and choreography complexity, while maintaining competitive scores across other metrics. Table 2 demonstrates our superiority in video quality metrics, particularly in imaging quality and overall consistency compared to MM-Diffusion [39] and Mochi [44]. In Table 3, MusicInfuser shows improved prompt alignment over the baseline Mochi model, with significant gains in creative interpretation and overall satisfaction.
Ablation Studies
Our ablation study (Table 4) systematically evaluates the contribution of each component in our framework. The full model achieves the highest overall score, with the layer adaptability selection providing the most significant contribution. HR-LoRA contributes substantially to movement realism, while our Beta-Uniform scheduling improves body representation. The naive feature addition baseline, where instead of using the ZICA adapter we simply spatially expand the audio feature and add it to the corresponding frame, performs worse than our approach in most metrics, confirming the effectiveness of our ZICA strategy. Additionally, in Table 3, we show the trade-off between style capture and creative interpretation of the prompt depending on the base prompt ratio of 0% and 100%.
Diversity of Results
By varying the random seed while keeping the prompt and music constant, our model generates diverse choreographies as shown in Fig. 8, demonstrating that it does not simply memorize specific dance routines for particular music tracks but instead learns a rich mapping between music and movement possibilities.
6 Conclusion
In this paper, we present MusicInfuser, a novel approach for generating dance videos synchronized with music by leveraging the rich choreographic knowledge embedded in pre-trained text-to-video diffusion models. Through our adaptation architecture and strategies, MusicInfuser successfully enables synchronized dance movements with musical inputs while preserving text-based control over style and scene elements. It achieves this without requiring expensive motion capture data, generalizes to novel music tracks, and supports the generation of diverse choreographies and group dance videos. By circumventing the prohibitive costs and restrictions of conventional approaches, MusicInfuser opens new possibilities for creative music-driven dance video synthesis.
Acknowledgments
Supplementary Material
Appendix A Video Results
We present the flattened video results along the time axis and the corresponding spectrograms in the main paper. However, our frame sampling rate does not exceed the Nyquist frequency for the general musical beat, causing the movement to appear slower. Therefore, we encourage readers to visit the project page to view the video results.
Appendix B Beta Distribution
Fig. 9 shows the evolution of graphs of Beta distributions from $\text{Beta}(1,3)$ to $\text{Beta}(1,1)$.
Figure 9: Beta distributions.
Appendix C Difficulty Control
@@ Appendix D Limitations
Appendix E Comparison with the Base Model
We show a comparison with Mochi [44] in Fig. 11. Note that Mochi is not able to hear the music.
Figure 11: MusicInfuser infuses listening capability into the text-to-video model (Mochi [44]), while preserving the prompt adherence and improving overall consistency and realism (frames 2 and 5 of the top example, and frames 2–4 of the bottom example).
Appendix F Ablation Study
We present qualitative results of our ablation study in Fig. 12 and Fig. 13. Our full model successfully generates consistent body shapes that align with the music, while preserving prior knowledge without introducing significant artifacts.
@@ Appendix G Layer Adaptability
The imaging and aesthetic quality of the base model [44] is presented in Fig. 14. This is analyzed with STG [26], an inference-time technique, and the score is calculated with VBench [25]. Based on the imaging quality, we choose the top 16 out of 48 layers in terms of image quality to select the layers.
@@
Appendix H Test Music Tracks
@@
Appendix I In-the-Wild Music Tracks
We use in-the-wild music tracks to demonstrate the generalizability of our method. In the main paper, we use SUNO AI to generate the music tracks, which produces music for non-commercial use. In addition, in Fig. 15, we show the results generated with three distinct music tracks from TikTok videos.
Figure 15: Videos generated with three distinct in-the-wild music tracks.
@@ Appendix I In-the-Wild Music Tracks
As mentioned in our main paper, we use a proper prompt format and base prompt for AIST [46]. The full list is shown in Table 6. Note that since we use VideoChat2 [31] to label YouTube videos, we have only the base prompt for the dataset. We also provide a predefined set of prompts in Table 7 that is used to generate samples for the evaluation, which ultimately results in $10\times 10=100$ videos for evaluation for each model configuration. The system prompts for VideoLLaMA [14] used for evaluation are in Table 8.

### v2

@@ 4.3 Zero-Initialized Adaptation Modules
4.4 Utilizing In-the-Wild Data
Training exclusively on datasets in highly constrained settings [45, 32] can lead to reduced generalizability and model degradation when confronted with diverse real-world scenarios. Therefore, we use a mixture of in-the-wild data and the constrained datasets. These videos introduce diversity in terms of camera trajectories, lighting conditions, performance environments, and dance styles. The inclusion of in-the-wild data serves as regularization, preventing overfitting to specific dance patterns or environmental settings. Details are provided in the supplementary material.
4.5 Prompt Diversification
In addition, we randomly replace a small portion of detailed captions with basic, simple captions. This allows the adapter network to learn how to respond to music without relying on the text, effectively reducing the model’s dependence on textual cues and encouraging it to develop stronger associations between musical features while still maintaining prompt adherence. This trade-off between style capture and music interpretation of the prompt is reflected in Table 4. The prompt templates used for diversification and replacement are provided in the supplementary material.
5 Experiments
5.1 Implementation Details
Model Details
We train our model on a single NVIDIA A100 GPU (except for experiments requiring more capacity) for 4,000 steps with a learning rate of $1\text{e}{-4}$, which takes roughly 20 hours to complete. Our LoRA uses rank 64, providing sufficient capacity to capture complex dance movements while maintaining parameter efficiency. For Beta-Uniform scheduling, we set the initial $\beta=3$ with exponential decay toward $\beta=1$. We use Mochi [43] as our base model with a classifier-free guidance scale of $\gamma_{\text{cfg}}=6.0$ during inference and employ Wav2Vec 2.0 [6] as the audio encoder, while using a shallow MLP followed by downsampling to match the temporal dimension of the audio tokens for the audio projector.
Dataset
The AIST dataset [45] includes 13,940 videos with 60 musical pieces, 10 dance genres, and 35 dancers. We extract 2,378 clips and divide the training and test sets with non-overlapping music tracks, following AIST++ [32]. We randomly sample approximately 2.5-second clips from full sequences for training. As mentioned in Sec. 4.4, we supplement AIST with 15,799 in-the-wild dance video clips from 4 YouTube playlists containing over 3.7k videos across various dance styles and settings. These clips are mixed with the AIST data at a 1:1 ratio during training, creating a balanced dataset that combines AIST’s controlled studio environment with diverse real-world performances.
Quantitative Metrics
Evaluating generated content automatically presents challenges. Inspired by VBench’s use of Visual-Language Models for text-to-video assessment [31], we propose a novel metric using VideoLLaMA 2 [14] and Qwen3-Omni [46], which process both video and audio inputs. We formulate targeted queries to assess three components: dance quality (style alignment, beat alignment, body representation, movement realism, choreography complexity), video quality (imaging quality, aesthetic quality, overall consistency), and prompt alignment (style capture, creative interpretation, satisfaction). See the supplementary material for exact prompts and methods. Tables 2 and 3 also display results on AIST test data, where the ground truth data outperforms generated content in metrics like beat alignment and movement realism. This validates that our metric correctly assigns higher scores to ground truth data, which should represent the upper bound for these metrics, and demonstrates its reliability alongside correlation with human evaluation results shown in Fig. 9.
5.2 Experimental Results
Music- and Text-Driven Dance Video Generation
Fig. 1 showcases the model’s ability to combine textual control with musical synchronization. The generated videos successfully incorporate scene contexts (restaurant kitchen, beach at sunset) and dancer attributes (wearing a leather jacket, chef’s uniform) as specified in the prompts, while simultaneously aligning the choreographic style with the musical input. Figs. 3 and 4 demonstrates that our model is capable of generating unseen subjects or rare settings.
Music Responsiveness
@@ 5.2 Experimental Results
Generalization to In-the-Wild Music and Longer Videos
To evaluate generalization beyond the AIST music distribution, we test our model with music tracks generated by SUNO AI. Fig. 6 shows successful generation for these unseen music categories, confirming the model’s ability to map novel audio patterns to appropriate dance movements. In addition, Fig. 6 shows longer video generation results with the same setting but with multiple times as many frames as the videos we used for training, up to 9 seconds.
Baseline Comparison
We present several baselines for our layer adaptation in Table 1. Adapting the layers with cross-attention that we selected based on the layer adaptability criterion in Sec. 4.1 significantly outperforms the strategy of selecting evenly distributed layers, first layers, middle layers, and last layers, and even outperforms adapting all layers in the video diffusion model. This demonstrates that our positive influence function for layer selection is crucial for high-performance adaptation.
Tables 23 present our quantitative results compared against prior work [38, 43]. For dance quality (Table 2), our method outperforms previous approaches in style alignment, beat alignment, movement realism, and choreography complexity, while maintaining competitive scores across other metrics. Table 3 demonstrates our superiority in video quality metrics, particularly in imaging quality and overall consistency compared to MM-Diffusion [38] and Mochi [43]. In Table 4, MusicInfuser shows improved creative interpretation and overall satisfaction over the baseline Mochi model. For qualitative comparisons with prior work, we refer readers to the supplementary material.
Human Evaluation
We conduct human evaluation to validate MusicInfuser’s performance and examine the correlation between Video-LLM-based quantitative assessment (Tables 2 and 3) and human judgments. Fig. 9 presents the results of our human evaluation study, where we assess generated videos across multiple dimensions including video quality, music-dance alignment, motion realism, and choreography complexity. The human evaluation demonstrates that our approach consistently outperforms previous work, with evaluators particularly noting improvements in video quality and movement naturalness. The details are provided in the supplementary material.
Figure 9: Human evaluation.
Ablation Studies
In Table 1, we evaluate the contribution of the components in our framework. Using Beta-Uniform scheduling improves body representation and movement realism. The naive feature addition baseline, where instead of using the ZICA adapter we simply spatially expand the audio feature and add it to the corresponding frame, similar to ControlNet [49], performs worse than our approach in most metrics, confirming the effectiveness of our ZICA strategy. Not zero-initializing the cross-attention layers results in a remarkable drop in the video quality metric. Additionally, in Table 4, we show the trade-off between style capture and creative interpretation of the prompt depending on the base prompt ratio, meaning how frequently we replaced the prompt with the basic prompt. More ablation studies and analysis are in the supplementary material.
Diversity of Results
By varying the random seed while keeping the prompt and music constant, our model generates diverse choreographies as shown in Fig. 8, demonstrating that it does not simply memorize specific dance routines for particular music tracks but instead is able to generate diverse dance sequences.
6 Conclusion
In this paper, we present MusicInfuser, a novel approach for generating dance videos synchronized with music by leveraging the rich choreographic knowledge embedded in pre-trained text-to-video diffusion models. Through our adaptation architecture and strategies, MusicInfuser successfully enables synchronized dance movements with musical inputs while preserving text-based control over style and scene elements. It achieves this without requiring expensive motion capture data, generalizes to novel music tracks and subjects, and supports the generation of diverse choreographies.
Acknowledgments
Supplementary Material
Figure 10: Comparison of audio-driven generation with MM-Diffusion [38]. Our method produces fewer artifacts (shown in the first and third rows), while generating more realistic dance videos with more natural movements (first row) and more dynamic motion (second and third rows). Note that we use the same music track for each row, and the spectrogram is stretched for $M$-Diffusion since we generate longer videos. For our method, we use the fixed caption “a professional dancer dancing …” across all music tracks.
Figure 11: MusicInfuser infuses listening capability into the text-to-video model (Mochi [43]), while preserving the prompt adherence and improving overall consistency and realism.
Appendix A Video Results
We present the flattened video results along the time axis and the corresponding spectrograms in the main paper. However, our frame sampling rate does not exceed the Nyquist frequency for the general musical beat, causing the movement to appear slower. Therefore, we encourage readers to view the supplementary video.
Layer Selection Strategy Style Beat Body Movement Choreography Dance Quality Imaging Aesthetic Overall Video Quality Overall
Layer Adaptability 7.56 8.89 7.16 8.24 7.90 7.95 9.60 7.87 9.39 8.95 8.33
Evenly Distributed Layers 7.31 8.81 7.28 7.70 7.96 7.81 9.33 7.78 9.04 8.72 8.15
First Layers 7.25 8.82 6.86 7.37 8.05 7.67 9.66 7.91 9.27 8.95 8.15
Middle Layers 7.91 8.87 6.74 7.83 7.98 7.86 9.21 7.97 9.20 8.79 8.21
Last Layers 7.52 8.81 7.01 7.47 8.00 7.76 9.45 7.73 9.14 8.77 8.14
All Layers 7.49 8.53 6.72 8.16 7.85 7.75 9.33 7.99 9.11 8.81 8.15
Table 5: Evaluation of layer selection strategies using VideoLLaMA 2 [14].
@@ Appendix A Video Results
Table 6: Ablation study. Feature addition denotes that we spatially expand the audio feature and add it to the corresponding frame. We use VideoLLaMA 2 [14] for the evaluation.
Appendix B Dance Difficulty Control
@@
Appendix C Human Evaluation Protocol
For each test music track [32], we conducted fully anonymized A/B testing. We asked 33 participants to evaluate the video quality, music-dance alignment, motion realism, and choreography complexity. The following are examples of the questionnaire items:
  1. 1.
    Which video has higher visual quality?
  2. 2.
    Which video’s dance aligns better with the music?
  3. 3.
    Which video’s motion is more realistic?
  4. 4.
    Which video’s dance is more complex?
@@ Appendix D Limitations
Appendix E Additional Qualitative Analysis
Fig. 10 presents a side-by-side comparison with MM-Diffusion [38]. Unlike MM-Diffusion, which generates shorter videos with limited style control, MusicInfuser produces longer sequences with both musical synchronization and prompt-based style control while improving the overall consistency of the video and reducing artifacts. We show a comparison with Mochi [43] in Fig. 11. Compared to Mochi, MusicInfuser produces more consistent human forms, fewer visual artifacts, and more fluid, realistic movements. Our method adds music responsiveness while maintaining or improving video consistency. We also compare with EDGE [44], a state-of-the-art skeleton-based dance generation model, in the main paper.
We present qualitative results of our ablation study in Fig. 13 and Fig. 14. Our full model successfully generates consistent body shapes that align with the music while preserving prior knowledge without introducing significant artifacts.
Appendix F Additional Quantitative Analysis
Similar to the layer selection baselines and ablation studies in the main paper using Qwen3-Omni [46], we show evaluation using VideoLLaMA 2 [14] in Tables 5 and 6. The full model achieves the highest score. Using higher rank for LoRA contributes substantially to movement realism, while our Beta-Uniform scheduling improves body representation. The naive feature addition baseline, where instead of using the ZICA adapter we simply spatially expand the audio feature and add it to the corresponding frame, performs worse than our approach in most metrics, confirming the effectiveness of our ZICA strategy.
Additionally, in Table 4 in the main paper, we show the trade-off between style capture and creative interpretation of the prompt depending on the base prompt ratio, meaning how frequently we replaced the prompt with the basic prompt.
@@ Appendix G Layer Adaptability
The imaging and aesthetic quality of the base model [43] is presented in Fig. 15. This is analyzed with STG [26], an inference-time guidance method, and the score is calculated with VBench [25]. Based on the imaging quality, which is highly related to the structure and noisiness of the video samples, we select the top 16 out of 48 layers in terms of image quality.
@@
Appendix H Beta-Uniform Scheduling
The visualization of the Beta-Uniform scheduling strategy is shown in Fig. 16.
Figure 16: Beta distributions.
Appendix I Test Music Tracks
@@
As mentioned in our main paper, we use a proper prompt format and base prompt for AIST [45]. The full list is shown in Table 8. Note that since we use VideoChat2 [31] to label YouTube videos, we have only the base prompt for the dataset. We also provide a predefined set of prompts in Table 9 that is used to generate samples for the evaluation, which ultimately results in $10\times 10=100$ videos for evaluation for each model configuration. The system prompts for VideoLLaMA 2 [14] and Qwen3-Omni [46] used for evaluation are in Table 10.